What are the derivatives of all the trigonometric function like


sin x


cos x


tan x


?????

Grade:12

1 Answers

dev bhatnagar
15 years ago

Trgono metric Derivatives are 

  • d/dx(sin x) = cos x
  • d/dx(cos x) = -sin x
  • d/dx(tan x) = (sec x)2
  • d/dx(sec x) = (sec x)2(tan x)
  • d/dx(csc x) = -(csc x)2(cot x)
  • d/dx(cot x) = -(csc x)2